OpenStudy (anonymous):

I NEED HELP. does anybody know the answer to this? The graph of g(x) is f(x) translated to the left 8 units and up 2 units. What is the function rule for g(x) given f(x) = x²?

OpenStudy (swissgirl):

\(g(x)=(x+8)^2+2\)

OpenStudy (swissgirl):

\(f(x)=(x-s)^2+t\) S= Horizontal translation. If the graph moved right or left T=Vertical translation. If the graph moved up or down

OpenStudy (anonymous):

ah i was right thank you! That was my answer i just wasnt sure!